Scrum of scrums/2013-11-12
Notable action items
NB these are action points that came up during the meeting; there are other blockers and important bits otherwise buried in the notes.
- Language team needs help from Ops setting up a dev environment for testing replag (https://mingle.corp.wikimedia.org/projects/scrum_of_scrums/cards/30)
- Operations blocked by W0 on varnish conf changes (https://mingle.corp.wikimedia.org/projects/scrum_of_scrums/cards/17)
- Mobile apps blocked on Wikidata for enhanced search (https://mingle.corp.wikimedia.org/projects/scrum_of_scrums/cards/42)
- Coordination needed between Analytics/W0 folks X-Forwarded-For header stuff (https://mingle.corp.wikimedia.org/projects/scrum_of_scrums/cards/20)
- Platform blocked by Ops on article import from production to betalabs (https://mingle.corp.wikimedia.org/projects/scrum_of_scrums/cards/37)
- Mobile Web still blocked by Growth on Getting Started API (https://mingle.corp.wikimedia.org/projects/scrum_of_scrums/cards/11)
- W0 to discuss FirefoxOS with Mobile Apps (https://mingle.corp.wikimedia.org/projects/scrum_of_scrums/cards/43)
- QA to support getting Parsoid working on betalabs (https://mingle.corp.wikimedia.org/projects/scrum_of_scrums/cards/34)
- Mark Bergsma on holiday for three weeks starting 2013-11-18. You have been warned.
- Trevor going on paternity leave starting on 2013-11-15
- Entire VE team abroad, 2013-11-15, two weeks (india, london), back the night before Thanksgiving
- improving performance of universal language selector, may break some backwards compatibility, email sent to wikitech. Will affect Wikidata extension, deployed this Thursday.
- would like to have dev environment to test replication lag...? Waiting for feedback from ops. (https://mingle.corp.wikimedia.org/projects/scrum_of_scrums/cards/30)
- Measure Google Translations - on hold for now
Andrew Otto, Faidon Liambotis
- ULSFO and ESAMS rate limiting issues, folks are working on this.
- Site outage 2013-11-08, see engineering list. Please work on reducing error lines from fatal logs.
- Varnish-text is up and running! some issues to work on like elevated 503 responses.
- Squid->Varnish migration complete after 4 years. Squid to remain for another month or so as a rollback option \o/
- upload-lb.eqiad IP change for Zero will be completed before the end of the week
- deployment 'trebuchet' work: Ryan Lane is adding recursive submodule support, shadow reference repositories (pre-checked out reference clones during fetch stage) and MediaWiki localization generated on the targets, which uses the shadow reference support.
- Analytics row having PDU problems. Want to shuffle analytics boxes in different rows to reduce rack/row related SPOFs. This will probably pushback varnishkafka deployment :/
- Mark Bergsma going on vacation for 3 weeks starting next week
- Swift to be switched to eqiad on Thursday
- Blocked by W0 on varnish conf changes
- small patch for wikidata to get enhanced search, will follow up with deploy soon (https://mingle.corp.wikimedia.org/projects/scrum_of_scrums/cards/42)
- No significant changes since last time
- We're experiencing slave lag, had to rerun several reports.
- Took geowiki offline some weeks ago, but it is back in gerrit! Private data is back too, but not pubically accessible. Ask if you need it.
- X-forward-for handling, in discussiong with W0 folks, need to schedule time to implement solution. (https://mingle.corp.wikimedia.org/projects/scrum_of_scrums/cards/20)
- Integration (Antoine)
- see email asking for reviews for new Zuul/Gearman, Alex is reviewing and help Antoine deploy.
- DevOps Sprint
- ryan-deploy (what Andrew referred to) - trebuchet
- Devs See Ori's !logs on SAL :) upcoming mail to engineering@, everything that is fatal or error is now being logged. Fix these things!
- everything seems ok?
- Betalabs: no movement. Blocked due to ops resources, mostly time.
- Few UX changes.
- Currently have A/B test going on with Analytics.
- Running test on infobox viewing time.
- Trying to enable VisualEditor on Mobile Alpha site for tablet sized devices.
- Waiting to hear from Growth team on KeepGoing/GettingStarted
- XFF stuff is in progress,
- #17 W0 varnish configs: should be new patch out in next week or two for this.
- #2 ESI Headers: in holding pattern
- #3 Support SSL: Mark will be done with this by end of week, will be able to send new IPs to carriers.
- #38, XF-by review: have patch pending, waiting for review from ops. Allows to suppress banners for Opera mini.
- Meeting with Opera-Mini this afternoon, will be getting official IP info from them. Will loop in Analytics and Ops on anything relevant.
- Discussion on Firefox OS, will need to make decisions on features for apps, need to follow up with Brion (https://mingle.corp.wikimedia.org/projects/scrum_of_scrums/cards/43)
- Documented API at Parsoid#The_Parsoid_web_API
- Deployed with some web api library issues
- Enables private wiki support, implements cleaned-up DOM spec etc
- Need full-stack testing
- Public API soon at http://parsoidcache.svc.eqiad.wikimedia.org/ (https://gerrit.wikimedia.org/r/#/c/93527/)
- Rashomon testing going well so far: https://www.mediawiki.org/wiki/User:GWicke/Notes/Storage/Cassandra_testing
- Parsoid API,finding bugs, coordinating
- Parsoid on beta labs affecting, but Chris mentioned this already.